A body displaced a volume of 10ml and its mass is about 35g. Find it's density as well as it's specific gravity .


1
Expert's answer
2021-02-17T11:08:30-0500

(a)

"\rho_{body}=\dfrac{m_{body}}{V_{water}}=\dfrac{m_{body}}{V_{body}},""\rho_{body}=\dfrac{0.035\ kg}{10\cdot10^{-3} L\cdot\dfrac{0.001\ m^3}{1\ L}}=3500\ \dfrac{kg}{m^3}."

(b)

"SG=\dfrac{\rho_{body}}{\rho_{water}}=\dfrac{3500\ \dfrac{kg}{m^3}}{1000\ \dfrac{kg}{m^3}}=3.5."